# DiP: A Scalable, Energy-Efficient Systolic Array for Matrix Multiplication Acceleration

### v1

@@ III-A DiP Architecture
DiP is a scalable spatial architecture consisting of NxN PEs, designed to accelerate matrix multiplication computations, as shown in Fig. 2 (a). The input matrix moves diagonally across the PEs, passing from one row of PEs to the next. The boundary PEs are diagonally connected, such that the registered inputs of the leftmost PE column are connected to the inputs of the rightmost PE column in the subsequent row. The weight matrix is loaded vertically, and psums are accumulated vertically. Figure 2 (b) shows the architecture of each PE. The proposed PE uses a 2-stage pipelined MAC unit to perform multiply and accumulate operations. It employs four enabled registers for weight, input, multiplier output, and adder output. The weight and input registers are 8-bit, while the multiplier and adder registers are 16-bit. Control signals wshift, pe_en, mul_en, and adder_en manage the PE operations. Specifically, wshift enables the weight register, while pe_en enables the input register. mul_en and adder_en selectively enable their respective registers only during active computation cycles, reducing power consumption during inactive cycles. The wshift signal is shared across all PEs in the systolic array, whereas pe_en, mul_en, and adder_en are shared across each row.
@@ III-B DiP Dataflow
The proposed DiP architecture adopts a novel dataflow to control the movement of inputs, weights, and partial summations across the whole systolic array. DiP dataflow relies on two major upgrades compared to WS dataflow, the diagonal movement of input matrix, and the weight matrix permutation. Firstly, the input matrix moves diagonally across PE rows, as shown in Fig. 2 (a). Secondly, the proposed dataflow permutates the weights by shifting and rotating each column by its column index, as shown in Fig. 2 (c). The weights permutation is prepared offline on software level according to the permutation pseudo code in Fig. 3. For each column, it iterates over all rows, assigning each element in the permutated_matrix based on its row and column index. The permutation is done at software level or at run-time in memory at almost zero cost. This permutation eliminates the input and output synchronization FIFOs required by conventional WS systolic arrays. Moreover, it increases the PE utilization and throughput, and decreases the required chip area and latency.
Figure 4 shows a complete example for 3x3 DiP systolic array. It consists of three rows: a) Row-0: $PE_{00}$, $PE_{01}$, $PE_{02}$, b) Row-1: $PE_{10}$, $PE_{11}$, $PE_{12}$, c) Row-2: $PE_{20}$, $PE_{21}$, $PE_{22}$. The PE array is diagonally connected, as shown in Fig. 4 (a). The leftmost PE in each PE row is connected to the rightmost PE in the next PE row. The weight matrix is initially permutated to be prepared for weights loading. Each column is shifted and rotated by its column index, as shown in Fig. 4 (b). The weights are initially loaded, row by row, to the systolic array, as shown in Fig. 4 from Cycle -2 to Cycle 0. The loading of the last weight row and the loading of first input matrix are performed in parallel at Cycle 0. The input data is loaded in parallel to Row-0 including inputs for $PE_{00}$, $PE_{01}$, and $PE_{02}$. After accomplishing the required computations by Row-0, the input data is shifted diagonally from $PE_{00}$ to $PE_{12}$ and from $PE_{01}$ to $PE_{10}$ and from $PE_{02}$ to $PE_{11}$. Then after accomplishing the required computations by Row-1, the input data is shifted diagonally again from $PE_{10}$ to $PE_{22}$ and from $PE_{11}$ to $PE_{20}$ and from $PE_{12}$ to $PE_{21}$. The processing starts from Cycle 1 to Cycle 5, while the first output row becomes ready at Cycle 3, and the last output row becomes ready at Cycle 5. The processing goes as follows:
@@
  • Cycle -1: The last row of weights matrix (c, d, h) is shifted to the second PE row, and new weights row (b, f, g) is loaded to the first PE row.
  • Cycle 0: The last row of the weight matrix (c, d, h) is shifted to the last PE row, the weights in the first PE row (b, f, g) are shifted to the second PE row, and new weights row (a, e, i) is loaded to the first PE row. To save one cycle, the first input matrix row (1, 2, 3) is loaded to the first PE row, simultaneously.
@@
  • Cycle 3: The first PE row shift the partial summations (7a, 8e, 9i) to the second row, and the input matrix row (7, 8, 9) is permutated to (8, 9, 7) and loaded to the second PE row. Similarly, The second PE row shift the partial summations (4a+5b, 5e+6f, 6i+4g) to the second row, and the input matrix row (5, 6, 4) is permutated to (6, 4, 5) and loaded to the third PE row. In addition, The third PE row shift the first output row (1a+2b+3c, 2e+3f+1d, 3i+1g+2h).
  • Cycle 4: The first PE row becomes idle, unless more input rows are loaded, and the second PE row shift the partial summations (7a+8b, 8e+9f, 9i+7g) to the third row, and the input matrix row (8, 9, 7) is permutated to (9, 7, 8) and loaded to the third PE row. In addition, the third PE row shift the second output row (4a+5b+6c, 5e+6f+4d, 6i+4g+5h).
  • Cycle 5: The first and second PE rows are idle, and the third PE row shift the third output row (7a+8b+9c, 8e+9f+7d, 9i+7g+8h).
Meanwhile, more new inputs may loaded if the input matrix is larger, and the processing continues till the end of the workload.
Figure 5: (a) Latency (per single tile processing) for both WS and DiP systolic arrays. The grey curve indicates the percentage of saved latency of DiP over WS, (b) Throughput, measured in operations per cycle (OPS/Cycle), for WS and DiP, with the grey curve showing the throughput improvement percentage of DiP compared to WS, (c) The number of used registers for DiP compared to WS, normalized to 8-bit (baseline datawidth). The grey curve represents the percentage of saved registers, and (d) TFPU represented the time required to activate all PEs, with the grey curve representing the TFPU improvement percentage. WS and DiP are shown in blue and black, respectively.
@@ III-C DiP Analytical Model
The analytical models for DiP are studied for latency and throughput. Regarding DiP latency, DiP systolic array consumes 2N+S-2 cycles for processing, where N is the number of rows/columns per DiP systolic array, and S is the number of pipelined stages per MAC unit. As a result, it takes 2N-1 cycles for 1-stage pipelined PE, and 2N cycles for 2-stage pipelined PE, as shown in (5). The throughput is calculated as the number of operations (multiplications and additions) divided by the latency, as shown in (6). TFPU is calculated, as shown in (7), where it takes $N$ cycles to reach full PEs utilization, outperforming WS by $N-1$ cycles. For the FIFO overhead, DiP systolic array eliminates the FIFOs overhead passing the whole input row in parallel without using any input synchronization FIFOs. Correspondingly, the output is generated row by row without any need of output synchronization FIFOs.
@@ III-D DiP/WSSA Analytical Comparison
The scalability of systolic arrays is important to meet the acceleration requirements. The proposed systolic array is gradually scaled up from 3x3 to 64x64 with sizes (3x3, 4x4, 8x8, 16x16, 32x32, 64x64). An analytical comparison of DiP and WS is conducted, evaluating throughput, latency, register savings, and TFPU across different systolic array sizes. Figure 5 (a) shows the latency for DiP compared to WS, with the percentage of latency savings calculated as the difference between WS and DiP latencies, divided by WS latency. The saved percentage starts at 28% for a 3x3 systolic array and reaches 33% for a 64x64 systolic array.
@@
Moreover, Fig. 5 (c) shows the percentage of saved registers as another design improvement of the proposed design compared to WS systolic array. By eliminating input/output FIFOs leading to percentage of saved registers reach up to 20% for 64x64 systolic array. The saved registers is calculated as the difference between WS and DiP used registers, divided by the number of registers used by WS. The registers of WS systolic array is distributed between input synchronization FIFOs, Output synchronization FIFOs, and internal PE registers. In contrast, the proposed diagonal input systolic array relies solely on internal PE registers, eliminating the need for input/output FIFOs. The represented number of registers are normalized to 8-bit as the baseline bandwidth.
TFPU calculates the required number of cycles to reach full utilization of PEs . This metric shows the overhead when the inputs are initially loaded, particularly for large matrix-matrix multiplication. Figure. 5 (d) shows TFPU for WS and DiP systolic arrays. The proposed DiP rapidly utilizes all PEs row by row, whereas WS gradually activates PEs in a diagonal pattern, starting from the top-left and moving to the bottom-right. Consequently, DiP outperforms WS, requiring only almost half the time of WS to fully utilize the entire systolic array.
IV Evaulation & Results
@@
TABLE I: Comparison of Area, Power Consumption, and Saving Percentages for Different WS/DiP Sizes using Commercial 22nm Technology at Frequency of 1GHz
Size Area Saved Area (%) Power (mW) Saved Power
Consumption (%)
@@
4x4 5,178  ($\mu$m2) 4,872  ($\mu$m2) 5.91 4.168 3.582 14.06
8x8 18,703 ($\mu$m2) 17,376 ($\mu$m2) 7.10 16.2 13.72 15.31
16x16 71,204 ($\mu$m2) 65,421 ($\mu$m2) 8.12 64.28 53.63 16.57
32x32 0.275 (mm2) 0.253 (mm2) 7.97 264.2 211.5 19.95
64x64 1.085 (mm2) 1.012 (mm2) 6.73 1041 857.8 17.60
TABLE II: Improvements in Throughput, Power Consumption, Area, and Overall Performance of DiP Compared to WS at Different Sizes
Size Throughput
Improvement ($\times$)
Power Consumption
Improvement ($\times$)
Area
Improvement ($\times$)
Overall
Improvement* ($\times$)
4x4 1.38 1.16 1.06 1.70
8x8 1.44 1.18 1.08 1.84
16x16 1.47 1.20 1.09 1.93
32x32 1.48 1.25 1.09 2.02
64x64 1.49 1.21 1.07 1.93
*Overall improvement represents energy efficiency per area
@@ IV-A Hardware Design Space Exploration
A hardware design space exploration is developed for DiP and WS at different sizes. Both designs are scaled from 4x4 to 64x64 with variants (4x4, 8x8, 16x16, 32x32, 64x64). A parameterized HDL design using Verilog is developed. Then, all designs are implemented from synthesize to GDSII using commercial 22nm technology at frequency of 1GHz. Table I shows a comparison between WS and DiP at different sizes in terms of area and power consumption. Additionally, the saved area and power consumption are presented for each design. It is depicted that the saved area percentage reach up to 8.12%. For the power consumption, the saved percentage reaches up to 19.95%.
Table II presents improvements in throughput, power consumption, area, and overall improvement (energy efficiency per area) for different WS/DiP design space. DiP outperforms WS across all metrics, with overall performance from 1.7x to 2.02x. At a size of 32×32, DiP achieves 1.48× higher throughput than WS, 1.25× lower power consumption, and 1.09× smaller area footprint, resulting in a total improvement of 2.02×. Additionally, at size of 64x64, throughput is improved by 1.49x, power consumption is reduced by 1.21x, and area is decreased by 1.07x compared to WS, with overall improvement 1.93x.
@@ IV-B Transformers Benchmarking
In MHA, the input $X$ is first projected in Queries ($Q_{i}$), Keys ($K_{i}$), and Values ($V_{i}$) per each head $i$ using learned weight matrices, as shown in (8.1). The attention scores ($S_{i}$) for each head are computed by taking the scaled dot product of Queries ($Q_{i}$) and transposed Keys ($K_{i}$), followed by a softmax normalization, as described in (8.2). The attention scores ($S_{i}$) are multiplied with Values ($V_{i}$), producing the attention output $Attn_{i}$ for each head according to (8.3). The outputs from all heads are concatenated into a single matrix $\text{Attn}_{\text{concat}}$ (8.4), which is finally projected back to the model’s hidden dimension ($d_{\text{model}}$) using a learned output projection matrix $W^{O}$, as shown in (8.5). This process enables the model to capture information from multiple representation subspaces simultaneously, enhancing the model’s ability to represent complex sequences effectively.
$Q_{i}=XW_{i}^{Q},\quad K_{i}=XW_{i}^{K},\quad V_{i}=XW_{i}^{V}$ (8.1)
$S_{i}=\text{softmax}(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}})$ (8.2)
$\text{Attn}_{i}=S_{i}V_{i}$ (8.3)
$\text{Attn}_{\text{concat}}=\text{concat}(\text{Attn}_{1},\text{Attn}_{2},\dots,\text{Attn}_{h})$ (8.4)
$\text{MHA}=\text{Attn}_{\text{concat}}W^{O}$ (8.5)
Where $X$ be the input matrix. $W_{i}^{Q}$, $W_{i}^{K}$, and $W_{i}^{V}$ are the input projection weight matrices, and $Q_{i}$, $K_{i}$, and $V_{i}$ are the resulting query, key, and value matrices per head, respectively. $S_{i}$ is the score matrix, and $\text{Attn}_{i}$ is the attention output per head. After concatenating all attention outputs from $h$ heads, $\text{Attn}_{\text{concat}}$ is obtained. Finally, $W^{O}$ is the output projection weight matrix, and MHA is the final multi-head attention output.
FFN in transformers consists of two linear transformations with a non-linear activation function applied between them. The process begins with the first matrix multiplication, which projects the input $y$ into a higher-dimensional space using a weight matrix $W_{1}$ and adding bias $b_{1}$, as shown in (9.1). Next, a non-linear activation function, such as ReLU or GELU is applied. Finally, the second matrix multiplication applies the second linear transformation to the non-linearly transformed output, mapping it back to the original dimensionality, as shown in (9.2).
[…]
$\text{FFN}=\text{Z}\cdot W_{2}+b_{2}$ (9.2)
where $\mathbf{y}$ is the FFN input, $W_{1}$ is the first weight matrix, and $\mathbf{b}_{1}$ is the bias. $W_{2}$ is the second FFN weight matrix and $\mathbf{b}_{1}$ is the second bias. The output of the second transformation is the FNN result.
Table III presents the matrix multiplication dimensions for MHA and FFN operations within transformer models, highlighting the relationship between input sequence length ($l$), model size ($d_{\text{model}}$), head size ($d_{k}$), and FFN size ($d_{\text{FFN}}$). These dimensions provide insights into the computational requirements of transformer models, emphasizing how the sequence length and model size affect the matrix operations in both the attention mechanism and feed-forward network.
TABLE III: Matrix Multiplication Dimensions for MHA and FFN Workloads in terms of Sequence Length ($l$), Model Hidden Size ($d_{\text{model}}$), Head Size ($d_{k}$), FFN Size ($d_{\text{FFN}}$). Input matrices Sizes are $M\times N$ and $N\times K$, and Output Matrix is $M\times K$.
Stage Workload Dimensions
@@
MHA Input projections $Q_{i},K_{i},V_{i}$ $l\times d_{\text{model}}\times d_{k}$
Attention scores $Q_{i}K_{i}^{T}$ $l\times d_{k}\times l$
$\text{Attn}_{i}=S_{i}V_{i}$ $l\times l\times d_{k}$
$\text{Output projection Attn}_{\text{concat}}W^{O}$ $l\times d_{\text{model}}\times d_{\text{model}}$
FFN $W_{1}$ projection $l\times d_{\text{model}}\times d_{\text{FFN}}$
$W_{2}$ projection $l\times d_{\text{FFN}}\times d_{\text{model}}$

### v2

@@ III-A DiP Architecture
DiP is a scalable spatial architecture consisting of N×N PEs, designed to accelerate matrix multiplication computations, as shown in Fig. 2(a). The input matrix moves diagonally across the PEs, passing from one PE row to the next. The boundary PEs are diagonally connected, with the registered inputs of the leftmost PE column are connected to the inputs of the rightmost PE column in the subsequent row. The weight matrix is loaded vertically, and psums are accumulated vertically. Figure 2(b) shows the architecture of each PE.
The proposed PE is similar to WS PE, utilizing a 2-stage pipelined MAC unit and four enabled registers for input, weight, multiplier output and adder output. Control signals wshift, pe_en, mul_en, and adder_en manage the PE operations. Specifically, wshift enables the weight register, while pe_en enables the input register. mul_en and adder_en selectively enable their respective registers only during active computation cycles, reducing power consumption during inactive cycles. However in DiP, wshift, pe_en, mul_en, and adder_en are shared across each PE row.
@@ III-B DiP Dataflow
The proposed DiP architecture adopts a novel dataflow to control the movement of inputs, weights, and partial summations across the whole systolic array. DiP dataflow relies on two major upgrades compared to WS dataflow, the diagonal movement of input matrix, and the weight matrix permutation. Firstly, the input matrix moves diagonally across PE rows, as shown in Fig. 2(a). Secondly, the proposed dataflow permutates the weights by shifting and rotating each column by its column index, as shown in Fig. 2(c). The weights permutation is initially prepared on software level, as shown in Algorithm 1. For each column, it iterates over all rows, assigning each element in the permutated_matrix based on its row and column index. The permutation is performed at the software level when the second matrix is a weight matrix. Alternatively, it is executed at run-time by efficiently re-scheduling memory access across multi-bank memories with almost zero overhead. The proposed dataflow removes the need for input and output synchronization FIFOs typically required by conventional WS systolic arrays. Additionally, it enhances throughput and PE utilization while reducing chip area and latency.
Algorithm 1 : Matrix Permutation for DiP Dataflow
  Input: $matrix$
  Output: $permutated\_matrix$
  for $i\leftarrow 0$ to cols do
     for $j\leftarrow 0$ to rows do
        $permutated\_matrix[j][i]=matrix[(j+i)\%rows][i]$
     end for
  end for
Figure 3 shows a complete example for 3×3 DiP systolic array. It consists of three rows: a) Row-0: $PE_{00}$, $PE_{01}$, $PE_{02}$, b) Row-1: $PE_{10}$, $PE_{11}$, $PE_{12}$, c) Row-2: $PE_{20}$, $PE_{21}$, $PE_{22}$. The PE array is diagonally connected, as shown in Fig. 3(a). The leftmost PE in each PE row is connected to the rightmost PE in the next PE row. The weight matrix is initially permutated to be prepared for weights loading. Each column is shifted and rotated by its column index, as shown in Fig. 3(b). The weights are initially loaded, row by row, to the systolic array, as shown in Fig. 3 from Cycle -2 to Cycle 0. The loading of the last weight row and the first input matrix are performed in parallel at Cycle 0. The input data is loaded in parallel to Row-0 including inputs for $PE_{00}$, $PE_{01}$, and $PE_{02}$. After accomplishing the required computations by Row-0, the input data is shifted diagonally from $PE_{00}$ to $PE_{12}$ and from $PE_{01}$ to $PE_{10}$ and from $PE_{02}$ to $PE_{11}$. Then after accomplishing the required computations by Row-1, the input data is shifted diagonally again from $PE_{10}$ to $PE_{22}$ and from $PE_{11}$ to $PE_{20}$ and from $PE_{12}$ to $PE_{21}$. The processing starts from Cycle 1 to Cycle 5, while the first output row becomes ready at Cycle 3, and the last output row becomes ready at Cycle 5. The processing goes as follows:
@@
  • Cycle -1: The weights matrix row (c, d, h) is shifted to the second PE row, and new weights row (b, f, g) is loaded to the first PE row.
  • Cycle 0: The weight matrix row (c, d, h) is shifted to the last PE row, the weights in the first PE row (b, f, g) are shifted to the second PE row, and new weights row (a, e, i) and the first input matrix row (1, 2, 3) are loaded to the first PE row, simultaneously.
@@
  • Cycle 3: The first PE row shift the partial summations (7a, 8e, 9i) to the second row, and the input matrix row (7, 8, 9) is permutated to (8, 9, 7) and loaded to the second PE row. Similarly, The second PE row shift the partial summations (4a+5b, 5e+6f, 6i+4g) to the third row, and the input matrix row (5, 6, 4) is permutated to (6, 4, 5) and loaded to the third PE row. In addition, The first output row (1a+2b+3c, 2e+3f+1d, 3i+1g+2h) is shifted out.
  • Cycle 4: The first PE row becomes idle, unless more input rows are loaded, and the partial summations (7a+8b, 8e+9f, 9i+7g) are shifted from the second PE row to the third row. In addition, the input matrix row (8, 9, 7) is permutated to (9, 7, 8), and loaded to the third PE row. In addition, the second output row (4a+5b+6c, 5e+6f+4d, 6i+4g+5h) is shifted out.
  • Cycle 5: The first and second PE rows are idle, and the third output row (7a+8b+9c, 8e+9f+7d, 9i+7g+8h) is shifted out.
Meanwhile, additional associated input tiles to the loaded weight tile are scheduled for larger input matrix, and the processing continues till the end of the workload.
Figure 4: Analytical modeling results comparing WS and DiP in terms of latency, throughput, register usage, and TFPU. (a) Latency (per single tile processing) for WS and DiP systolic arrays, with the grey curve indicating the percentage of latency savings of DiP over WS. (b) Throughput, measured in operations per cycle (OPS/Cycle), for WS and DiP, with the grey curve showing the percentage throughput improvement of DiP compared to WS. (c) Register usage for DiP compared to WS, normalized to 8-bit (baseline datawidth), with the grey curve indicating the percentage of register savings. (d) TFPU, representing the time required to activate all PEs, with the grey curve showing the percentage TFPU improvement. WS and DiP are represented in blue and black, respectively.
@@ III-C DiP Analytical Model
The analytical models for DiP are studied for latency and throughput. Regarding DiP latency, DiP systolic array requires 2N+S-2 cycles for processing, where N is the number of rows/columns per DiP systolic array, and S is the number of pipelined stages per MAC unit. As a result, the latency is 2N-1 cycles for 1-stage pipelined PE, and 2N cycles for 2-stage pipelined PE, as shown in (5). The throughput is calculated as the number of operations (multiplications and additions) divided by the latency, as shown in (6). TFPU is calculated, as shown in (7), where it takes $N$ cycles to reach full PEs utilization, outperforming WS by $N-1$ cycles. Moreover, The registers overhead is eliminated by removing the input/output synchronization FIFOs.
@@ III-D DiP/WSSA Analytical Comparison
The scalability of systolic arrays is important to meet the acceleration requirements. The proposed systolic array is gradually scaled up from 3x3 to 64x64 with sizes (3x3, 4x4, 8x8, 16x16, 32x32, 64x64). An analytical comparison of DiP and WS is presented, evaluating throughput, latency, register savings, and TFPU across different systolic array sizes. Figure 4(a) shows the latency for DiP compared to WS, with the percentage of latency savings calculated as the difference between WS and DiP latencies, divided by WS latency. The saved percentage starts at 28% for a 3x3 systolic array and reaches 33% for a 64x64 systolic array.
@@
Moreover, Fig. 4(c) shows the percentage of saved registers as another design improvement of DiP compared to the WS systolic array. By eliminating input/output FIFOs, the percentage of saved registers reaches up to 20% for a 64×64 systolic array. The saved registers is calculated as the difference between WS and DiP used registers, divided by the number of registers used by WS. The registers of WS systolic array is distributed between input synchronization FIFOs, Output synchronization FIFOs, and internal PE registers. In contrast, the proposed diagonal input systolic array relies solely on internal PE registers, eliminating the need for input/output FIFOs. The represented number of registers are normalized to 8-bit as the baseline bandwidth.
TFPU calculates the required number of cycles to reach full utilization of PEs . This metric shows the overhead when the inputs are initially loaded, particularly for large matrix-matrix multiplication. Figure 4(d) shows TFPU for WS and DiP systolic arrays. The proposed DiP rapidly utilizes all PEs row by row, whereas WS gradually activates PEs in a diagonal pattern, starting from the top-left and moving to the bottom-right. Consequently, DiP outperforms WS, requiring only almost half the time of WS to fully utilize the entire systolic array.
IV Evaluation & Results
@@
TABLE I: Comparison of Area, Power Consumption, and Improvements in Throughput, Power Consumption, Area, and Overall Performance of DiP Compared to WS at Different Sizes using Commercial 22nm Technology at Frequency of 1GHz
Size Area (mm2) Area Improvement ($\times$) Power Consumption (mW) Power Consumption Improvement ($\times$) Throughput Improvement ($\times$) Overall Improvement* ($\times$)
@@
4×4 0.0052 0.0049 1.06 4.168 3.582 1.16 1.38 1.70
8×8 0.0187 0.0174 1.08 16.2 13.72 1.18 1.44 1.84
16×16 0.0712 0.0654 1.09 64.28 53.63 1.20 1.47 1.92
32×32 0.2750 0.253 1.09 264.2 211.5 1.25 1.48 2.02
64×64 1.085 1.012 1.07 1041 857.8 1.21 1.49 1.93
*Overall improvement represents energy efficiency per area.
@@ IV-A Hardware Design Space Exploration
A hardware design space exploration is developed for DiP and WS at different sizes. Both designs are scaled from 4x4 to 64×64 with variants (4×4, 8×8, 16×16, 32×32, 64×64). A parameterized HDL design using Verilog is developed. Then, all designs are implemented from synthesize to GDSII using commercial 22nm technology at frequency of 1 GHz. Table I shows a comparison between WS and DiP at different sizes in terms of area and power consumption. Additionally, the improvements in throughput, power consumption, area, and overall improvement (energy efficiency per area) for different WS/DiP design space. DiP outperforms WS across all metrics, with overall performance from 1.7× to 2.02×. At a size of 32×32, DiP achieves 1.48× higher throughput than WS, 1.25× lower power consumption, and 1.09× smaller area footprint, resulting in a total improvement of 2.02×. Additionally, at size of 64×64, the throughput is improved by 1.49×, power consumption is reduced by 1.21×, and area is decreased by 1.07× compared to WS, with overall improvement 1.93×.
@@ IV-B Transformers Benchmarking
Table II presents the matrix multiplication workloads and their dimensions for MHA and FFN operations within transformer models, highlighting the relationship between input sequence length ($l$), model size ($d_{\text{model}}$), head size ($d_{k}$), and FFN size ($d_{\text{FFN}}$). The table offers insights into the computational requirements of transformer models, emphasizing how the sequence length and model size impact the matrix operations in both the attention mechanism and the feed-forward network. In MHA, the input ($X$) is first projected in Queries ($Q_{i}$), Keys ($K_{i}$), and Values ($V_{i}$) per each head $i$ using learned weight matrices. The attention scores ($S_{i}$) for each head are computed by taking the scaled dot product of Queries ($Q_{i}$) and transposed Keys ($K_{i}$), followed by a softmax normalization. The attention scores ($S_{i}$) are multiplied with Values ($V_{i}$), producing the attention output ($Attn_{i}$) for each head. The outputs from all heads are concatenated into a single matrix $\text{Attn}_{\text{concat}}$, which is finally projected back to the model’s hidden dimension ($d_{\text{model}}$) using a learned output projection matrix ($W^{O}$). This process enables the model to capture information from multiple representation subspaces simultaneously, enhancing the model’s ability to represent complex sequences effectively. FFN in transformers consists of two linear transformations with a non-linear activation function applied between them. The process begins with the first matrix multiplication, which projects the input $Y$ into a higher-dimensional space using a weight matrix $W_{1}$ and adding bias $b_{1}$, as shown in Table II. Next, a non-linear activation function, such as ReLU or GELU is applied. Finally, the second matrix multiplication applies the second linear transformation to the non-linearly transformed output, mapping it back to the original dimensionality.
TABLE II: Matrix Multiplication Dimensions for MHA and FFN Workloads in terms of Sequence Length ($l$), Model Hidden Size ($d_{\text{model}}$), Head Size ($d_{k}$), FFN Size ($d_{\text{FFN}}$). Input matrices Sizes are $M\times N$ and $N\times K$, while Output Matrix is $M\times K$.
Workload Dimensions
@@
Input Projections, $[Q_{i},K_{i},V_{i}]=X\cdot[W^{Q}_{i},W^{K}_{i},W^{V}_{i}]$ $l\times d_{\text{model}}\times d_{k}$
$A_{i}=Q_{i}\cdot K_{i}^{T}$ $l\times d_{k}\times l$
Attention Scores, $S_{i}=\text{Softmax}(A_{i}/\sqrt{d_{k}})$ $l\times d_{k}\times l$
Attention, $\text{Attn}_{i}=S_{i}\cdot V_{i}$ $l\times l\times d_{k}$
Output Projection, $Y=\text{Concat}(\text{Attn}_{1},\text{Attn}_{2},\dots,\text{Attn}_{h})\cdot W^{O}$ $l\times d_{\text{model}}\times d_{\text{model}}$
FFN1 Projection, $Z=\text{Non-Linear}(Y\cdot W_{1}+b_{1})$ $l\times d_{\text{model}}\times d_{\text{FFN}}$
FFN2 Projection, $F=Z\cdot W_{2}+b_{2}$ $l\times d_{\text{FFN}}\times d_{\text{model}}$
Figure 5: Evaluation of DiP and TPU-like architecture at size 64×64 with MHA and FFN transformer workloads. The evaluation includes actual energy consumption (a, b) and latency (c, d) across various workloads dimensions of matrix multiplication (M-N-K).